# Multi-Station WiFi CSI Sensing Framework Robust to Station-wise Feature Missingness and Limited Labeled Data

### 12 Mar 2026 v1

@@ III Proposed Method
Figure 2: Illustration of labeled and unlabeled sample generation from asynchronous CSI and label streams. Both labeled and unlabeled samples are constructed by applying a fixed-length window around a reference timestamps. Due to asynchronous sensing and heterogeneous transmission intervals, some windows may contain no CSI from particular stations, resulting in station-wise feature missingness.
@@ III-B Framework Overview
The proposed framework consists of two steps: pre-training with an unlabeled dataset and downstream model training with a labeled dataset. Here, we briefly describe the training procedure, with details deferred to SubSection III-C, III-D. In the pre-training, we aim to learn a station-missingness-invariant global representation from unlabeled CSI data. To explicitly model station-wise missingness during pre-training, we employ cross-modal self-supervised learning (CroSSL) [5]. Each station-specific input $X_{i}^{(d)}$ is encoded by $E^{(d)}$ to obtain an intermediate embedding $Q_{i}^{(d)}$}. To simulate station-wise feature missingness, two artificial masking sets $\mathcal{M}_{i}^{\mathrm{mask}},\mathcal{M}_{i}^{\mathrm{mask}^{\prime}}\subseteq\{1,\ldots,N_{d}\}$ is sampled, and the embeddings corresponding to stations in $\mathcal{M}_{i}^{\mathrm{mask}},\mathcal{M}_{i}^{\mathrm{mask}^{\prime}}$ are replaced with a masked value $M_{\mathrm{emb}}$. The masked embeddings are then fused by a shared aggregator $A(\cdot)$ to produce two global embeddings, $Z_{i}$ and $Z_{i}^{\prime}$. A self-supervised loss is imposed between $Z_{i}$ and $Z_{i}^{\prime}$, encouraging consistent global representations under different station-wise missingness patterns. By enforcing agreement between representations derived from different subsets of available stations, the model is discouraged from relying on any specific station and instead learns station-missingness-invariant global representations. As a result, the feature extractor consisting of $E^{(d)}(\cdot)$ and $A(\cdot)$ learns representations invariant to station-wise feature missingness while leveraging unlabeled data.
Next, the downstream model is trained with limited labeled data. While CroSSL encourages the feature extractor to produce representations invariant to missing stations, a mismatch can still arise if the downstream model is trained only on fully observed multi-station inputs. To address this issue, we introduce Station-wise Masking Augmentation (SMA) during downstream model training. Given a labeled multi-station input $X_{i}$ with ground-truth $Y_{i}$, we first sample a station-wise masking set $\mathcal{M}_{i}^{\mathrm{mask}}\subseteq\{1,\ldots,N_{d}\}$. The inputs corresponding to stations in $\mathcal{M}_{i}^{\mathrm{mask}}$ are then replaced with a masked value $M_{\mathrm{inp}}$. The masked input is processed by the same station-specific encoders and aggregator as in pre-training to obtain a global embedding $Z_{i}$, which is then fed into a downstream model $D(\cdot)$ to produce a prediction $\hat{Y}_{i}$. The downstream model is trained by minimizing a supervised loss between $\hat{Y}_{i}$ and the ground-truth $Y_{i}$. By explicitly exposing the downstream model to station-wise missingness during training, SMA ensures that the entire sensing models remains robust when stations are unavailable at inference time.
@@ III-D Downstream Model Training with SMA
$\mathbb{P}(d\in\mathcal{M}_{i})=p_{\mathrm{mask}}.$ (19)
@@
IV Evaluation using office like environment dataset
To verify the effectiveness of the proposed method, we conducted experiments using two datasets: office-like environment dataset and factory-like environment dataset. This section focuses on the evaluation using the office-like environment dataset. CSI and RGB images were collected while a single subject walked in a standard office environment. Using the collected data, we constructed a one-dimensional position estimation dataset and performed machine learning experiments to evaluate our proposed framework’s robustness to station-wise feature missingness and limited labeled data.
@@ IV-B Baseline
Fully Supervised Baselines: As fully supervised baselines, we introduce two methods trained solely on labeled data. The first is NaiveSupervised. For the office-like environment dataset, this baseline fed the CSI, concatenated along the station dimention, directory into the downstream model, bypassing any feature extractor. The second is an OutputEnsemble, in which $N_{d}$ separate downstream models are trained, and their outputs are averaged during inference. These methods serve as reference points for evaluating performance without pre-training.
@@ IV-D Robustness to Station-wise Feature Missingness
The proposed method consistently achieves the highest accuracy across a wide range of missingness levels. Even when more than half of the stations are unavailable, its performance degrades only marginally, demonstrating strong robustness to station-wise feature missingness. This indicates that the proposed method does not rely heavily on any particular station and retains its discriminative power even under substantial feature loss.
@@ IV-I Sensitivity Analysis of Masking Rate
Overall, robust performances are consistently observed with $p_{\mathrm{mask}}=0.5$. Additionally, when only one station is available at inference time, performance also drop with $p_{\mathrm{mask}}=0.1$ during both pre-training and downstream model training. This can be attributed to the mismatch between the low missingness rate assumed during training and the high level of missingness encountered during inference. These findings highlight that $p_{\mathrm{mask}}=0.5$ is a well-balanced choice for this dataset.
V Evaluation using factory like environment dataset
To further assess the robustness of proposed method, this section details an evaluation using factory-like environment dataset. The dataset was constructed by collecting CSI and RGB images as five subjects walked within a large, semi-enclosed indoor space. From this dataset, we formulated an image generation task.
@@ V-G Ablation Study of Pre-training and Augmentation Strategies
Table VII presents the results of different combinations of pre-training and data augmentation strategies on the factory-like environment dataset. Across all evaluated conditions, the proposed method and AutoFi × SMA consistently ranks among the top-performing approaches. In particular, the proposed method achieves the best performance in the office-like environment and maintains competitive performance in the factory-like environment, indicating that the combination of missingness-invariant pre-training and station-wise data augmentation is robust across multiple environments. These results suggest that the proposed method provides a reliable strategy for handling both station-wise feature missingness and limited labeled data under varying sensing conditions.
@@ VI Conclusion
In this study, we proposed a multi-station WiFi CSI sensing framework that addresses two practical challenges in real-world deployments: station-wise feature missingness and limited labeled data. The core idea is to learn representations that are invariant to station availability during self-supervised pre-training, and to propagate this robustness to downstream tasks through station-aware training. Specifically, we adapted CroSSL to the multi-station CSI setting to learn missingness-invariant representations from unlabeled data, and introduced Station-wise Masking Augmentation (SMA) to align downstream training with realistic station unavailability. Experimental results demonstrate that the joint design of CroSSL and SMA is essential for achieving robust performance under both station-wise feature missingness and label scarcity.
We evaluated the proposed framework on two real-world multi-station CSI datasets collected in office-like and factory-like environments. Across both datasets, the proposed method consistently outperformed conventional supervised, augmentation-based, and existing SSL baselines over a wide range of station-missingness levels and label usage ratios. These results indicate that explicitly incorporating station-wise missingness into both representation learning and downstream training provides a practical and effective foundation for deploying WiFi sensing systems in real environments, where incomplete inputs and limited labeled data are unavoidable.

### 24 Mar 2026 v2

@@ III Proposed Method
Figure 2: Illustration of labeled and unlabeled sample generation from asynchronous CSI and label streams. Both labeled and unlabeled samples are constructed by applying a fixed-length window around reference timestamps. Due to asynchronous sensing and heterogeneous transmission intervals, some windows may contain no CSI from particular stations, resulting in station-wise feature missingness.
@@ III-B Framework Overview
The proposed framework consists of two steps: pre-training with an unlabeled dataset and downstream model training with a labeled dataset. Here, we briefly describe the training procedure, with details deferred to SubSection III-C, III-D. In the pre-training, we aim to learn a station-missingness-invariant global representation from unlabeled CSI data. To explicitly model station-wise missingness during pre-training, we employ cross-modal self-supervised learning (CroSSL) [5]. Each station-specific input $X_{i}^{(d)}$ is encoded by $E^{(d)}$ to obtain an intermediate embedding $Q_{i}^{(d)}$}. To simulate station-wise feature missingness, two artificial masking sets $\mathcal{M}_{i}^{\mathrm{mask}},\mathcal{M}_{i}^{\mathrm{mask}^{\prime}}\subseteq\{1,\ldots,N_{d}\}$ are sampled, and the embeddings corresponding to stations in $\mathcal{M}_{i}^{\mathrm{mask}},\mathcal{M}_{i}^{\mathrm{mask}^{\prime}}$ are replaced with a masked value $M_{\mathrm{emb}}$. The masked embeddings are then fused by a shared aggregator $A(\cdot)$ to produce two global embeddings, $Z_{i}$ and $Z_{i}^{\prime}$. A self-supervised loss is imposed between $Z_{i}$ and $Z_{i}^{\prime}$, encouraging consistent global representations under different station-wise missingness patterns. By enforcing agreement between representations derived from different subsets of available stations, the model is discouraged from relying on any specific station and instead learns station-missingness-invariant global representations. As a result, the feature extractor consisting of $E^{(d)}(\cdot)$ and $A(\cdot)$ learns representations invariant to station-wise feature missingness while leveraging unlabeled data.
Next, the downstream model is trained with limited labeled data. While CroSSL encourages the feature extractor to produce representations invariant to missing stations, a mismatch can still arise if the downstream model is trained only on fully observed multi-station inputs. To address this issue, we introduce Station-wise Masking Augmentation (SMA) during downstream model training. Given a labeled multi-station input $X_{i}$ with ground-truth $Y_{i}$, we first sample a station-wise masking set $\mathcal{M}_{i}^{\mathrm{mask}}\subseteq\{1,\ldots,N_{d}\}$. The inputs corresponding to stations in $\mathcal{M}_{i}^{\mathrm{mask}}$ are then replaced with a masked value $M_{\mathrm{inp}}$. The masked input is processed by the same station-specific encoders and aggregator as in pre-training to obtain a global embedding $Z_{i}$, which is then fed into a downstream model $D(\cdot)$ to produce a prediction $\hat{Y}_{i}$. The downstream model is trained by minimizing a supervised loss between $\hat{Y}_{i}$ and the ground-truth $Y_{i}$. By explicitly exposing the downstream model to station-wise missingness during training, SMA ensures that the entire sensing model remains robust when stations are unavailable at inference time.
@@ III-D Downstream Model Training with SMA
$\mathbb{P}(d\in\mathcal{M}_{i}^{\mathrm{mask}})=p_{\mathrm{mask}}.$ (19)
@@
IV Evaluation using office-like environment dataset
To verify the effectiveness of the proposed method, we conducted experiments using two datasets: office-like environment dataset and factory-like environment dataset. This section focuses on the evaluation using the office-like environment dataset. CSI and RGB images were collected while a single subject walked in a standard office environment. Informed consent was obtained from the participant prior to data collection. Using the collected data, we constructed a one-dimensional position estimation dataset and performed machine learning experiments to evaluate our proposed framework’s robustness to station-wise feature missingness and limited labeled data.
@@ IV-B Baseline
Fully Supervised Baselines: As fully supervised baselines, we introduce two methods trained solely on labeled data. The first is NaiveSupervised. For the office-like environment dataset, this baseline fed the CSI, concatenated along the station dimension, directory into the downstream model, bypassing any feature extractor. The second is an OutputEnsemble, in which $N_{d}$ separate downstream models are trained, and their outputs are averaged during inference. These methods serve as reference points for evaluating performance without pre-training.
@@ IV-D Robustness to Station-wise Feature Missingness
The proposed method consistently maintains the lowest error across a wide range of missingness levels. Even when more than half of the stations are unavailable, its performance degrades only marginally, demonstrating strong robustness to station-wise feature missingness. This indicates that the proposed method does not rely heavily on any particular station and retains its discriminative power even under substantial feature loss.
@@ IV-I Sensitivity Analysis of Masking Rate
Overall, robust performances are consistently observed with $p_{\mathrm{mask}}=0.5$. Additionally, when only one station is available at inference time, performance also drops with $p_{\mathrm{mask}}=0.1$ during both pre-training and downstream model training. This can be attributed to the mismatch between the low missingness rate assumed during training and the high level of missingness encountered during inference. These findings highlight that $p_{\mathrm{mask}}=0.5$ is a well-balanced choice for this dataset.
V Evaluation using factory-like environment dataset
To further assess the robustness of proposed method, this section details an evaluation using factory-like environment dataset. The dataset was constructed by collecting CSI and RGB images as five subjects walked within a large, semi-enclosed indoor space. Informed consent was obtained from all participants prior to data collection. From this dataset, we formulated an image generation task.
@@ V-G Ablation Study of Pre-training and Augmentation Strategies
Table VII presents the results of different combinations of pre-training and data augmentation strategies on the factory-like environment dataset. Across all evaluated conditions, the proposed method and AutoFi × SMA consistently rank among the top-performing approaches. Although AutoFi × SMA achieves the lowest RMSE in most settings, the proposed method remains highly competitive across all conditions, consistently staying close to the best result. These results indicate that combining missingness-aware pre-training with station-wise augmentation is effective even for the more challenging image- generation task in the factory-like environment.
@@ VI Conclusion
In this study, we proposed a multi-station WiFi CSI sensing framework that addresses two practical challenges in real-world deployments: station-wise feature missingness and limited labeled data. The core idea is to learn representations that are invariant to station availability during self-supervised pre-training, and to propagate this robustness to downstream tasks through station-aware training. Specifically, we adapted CroSSL to the multi-station CSI setting to learn missingness-invariant representations from unlabeled data, and introduced Station-wise Masking Augmentation (SMA) to align downstream training with realistic station unavailability. We evaluated the proposed framework on two real-world multi-station CSI datasets collected in office-like and factory-like environments. Across both datasets, the proposed method consistently outperformed conventional supervised, augmentation-based, and existing SSL baselines over a wide range of station-missingness levels and label usage ratios. These results indicate that explicitly incorporating station-wise missingness into both representation learning and downstream training provides a practical and effective foundation for deploying WiFi sensing systems in real environments, where incomplete inputs and limited labeled data are unavoidable.
Future work includes expanding the variety of downstream tasks and model architectures to further demonstrate the generality of the proposed method. In addition, although this study assumed that signals from all stations contribute to the downstream task, this assumption may not hold in practice. Some stations may provide little useful information, such as mobile devices that continuously move. Therefore, it will be important to develop a mechanism for selecting the Stations to be used as features, including the prior exclusion of such non-contributory stations.